OpenStudy (anonymous):

John can paint a fence in 9 hours. Tim can paint the same fence in 11 hours. How much time would the two working together need to paint the same fence? (Round your answer to the nearest hundredth.)

OpenStudy (wolf1728):

There is a formula for that : 9 * 11 / ( 9+ 11) = 99 / 20

OpenStudy (mathmate):

John can finish 1/9 of the job in 1 hour. Tim can finish 1/11 of the job in 1 hour. Together they finish \( \frac{1}{9} + \frac{1}{11} \) in 1 hour. Add the fractions and take the reciprocal to get the time it takes to finish the job together. If you are good with memorizing formulas, memorize wolf1728's formula.

OpenStudy (mary.rojas):

There are some explanations on here as well http://answers.yahoo.com/question/index?qid=20090331115425AAMqmku

OpenStudy (wolf1728):

Basically, the formula is total time = (Rate 1 * Rate 2) / (Rate 1 + Rate 2)

OpenStudy (anonymous):

is the answer 4.95

OpenStudy (mathmate):

exactly!